OpenStudy (anonymous):

8x-(3x+6)=24 what is x?

OpenStudy (anonymous):

x=6

jimthompson5910 (jim_thompson5910):

8x-(3x+6)=24 8x-3x-6=24 5x-6=24 5x=24+6 5x=30 x=30/5 x=6

OpenStudy (anonymous):

8x-3x-6=24 5x=24+6 5x=30 x+6

OpenStudy (diyadiya):

5x-6=24 5x=30 x=6

OpenStudy (anonymous):

\[ \sqrt[3]{216} \]

OpenStudy (anonymous):

x=6*

jhonyy9 (jhonyy9):

5x=30 x=6